# Jamming is a first-order transition with quenched disorder in amorphous materials sheared by cyclic quasistatic deformations

### v1

@@ Conclusion.
In this study, we understand the nature of jamming transition through an ensemble approach analogous to the statistical mechanics in equilibrium systems Baule et al. (2018); Bi et al. (2015). Within such a framework, the order of the jamming transition is independent of the complex properties of jammed packings, including isostaticity, marginality and hyperuniformity (see (i-iii) in Introduction). As a first-order jamming transition, the finite-size scaling exponents measured in this study are naturally universal in 2D and 3D, without the need to introduce the upper critical dimension $d_{\rm u}=2$ Parisi et al. (2018), a concept relevant to continuous transitions only. It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction). The jamming transition appears to be a rare example with the simultaneous existence of first-order and second-order transitions between liquid and solid states. Conventionally, the two kinds of transitions only coexist in gas-liquid systems, but not in liquid-crystalline solid systems.
@@ Methods.
Shear protocol. Particles are randomly distributed at an initial volume fraction $\varphi_{0}=0.02$. The system is then rapidly quench compressed to the target $\varphi$. The CAQS is performed under the Lees-Edwards boundary conditions, with a fixed $\varphi$. We use a strain step $\delta\gamma=0.1$ to generate the phase diagram in Fig. 1a, and $\delta\gamma=0.01$ for other results. At each step, particle positions are shifted according to $x_{i}\rightarrow x_{i}+\delta\gamma y_{i}$, and then the system’s energy is minimized using the FIRE algorithm Bitzek et al. (2006). During a cycle, the strain is varied as $\gamma=0\rightarrow\gamma_{\rm max}\rightarrow 0$. We set $\gamma_{\rm max}=0.7$ in the main text and $\gamma_{\rm max}=1$ in the Extended Data for the 2D model, and $\gamma_{\rm max}=0.5$ for the 3D model (Extended Data). The number of cycles is represented by $t$ with unit oscillation period ($t=1$). In 2D, the maximum number of cycles is $t_{\rm max}=250$ for $0.839\leq\varphi\leq{0.849}$ and $t_{\rm max}=4000$ for $0.839{0.833}\leq\varphi<$ 0.839, while in 3D, $t_{\rm max}=144$ for $2000$, and $t_{\rm max}=50$ for $N=512$ and $1000$. We generate $N_{\rm s}=4000-12000$ independent samples at each $\varphi$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ configurations are collected to compute statistical quantities.
@@ Acknowledgements.
We warmly thank Hajime Yoshino for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.

### v2

@@ Conclusion.
In this study, we investigate the nature of the jamming transition through an ensemble approach analogous to the statistical mechanics in equilibrium systems Baule et al. (2018); Bi et al. (2015). Within such a framework, jamming is demonstrated to be a first-order transition with quenched disorder, in the quasistatic deformation limit where all fragile states are excluded. The order of the jamming transition is independent of the complex properties of jammed packings, including isostaticity, marginality and hyperuniformity (see (i-iii) in Introduction).
In previous rapid quench simulations, it is observed that the width $w$ of the jamming density distribution $P(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-\Omega}$ with $\Omega=0.55\pm 0.03$ in both 2D and 3D O’hern et al. (2003). This finite-size scaling depends on the total number of particles, $N$, rather than on the system length $\ell$, and the exponent $\Omega\approx 1/2$ is independent of the dimensionality within the numerical accuracy. An explanation, as suggested previously, is that jamming is a second-order transition with an upper critical dimension $d_{\rm u}=2$ Goodrich et al. (2012). Here we propose an alternative interpretation: jamming is a first-order transition with quenched disorder, and consequently the finite-size scaling is a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, independent of the dimensionality.
It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction). Conventionally, first- and second-order transitions coexist in gas-liquid systems, but not in liquid-crystalline solid systems. Because the liquid-crystalline solid transition is accompanied with spontaneous symmetry breaking, it cannot be a continuous type. However, there is no apparent spontaneous symmetry breaking during liquid-amorphous solid transitions (such as the jamming transition). Thus there is no fundamental reason to exclude the possibility of the coexistence of discontinuous and continuous transitions between a liquid and an amorphous solid. This study opens an avenue to unify first-order (this study) and second-order Olsson and Teitel (2007) jamming transitions. Another direction of the future work is to study the competition between the sample-to-sample fluctuations due to disorder and the fluctuations associated to the granular temperature Pöschel and Luding (2001); Luding et al. (2021), and examine the “thermal effects” on the nature of the first-order jamming transition. Similar studies on the thermal vestiges of the zero-temperature physics have been recently carried out in the driven RFIM at finite temperatures Yao and Jack (2023). It would be also interesting to examine if friction can alter the order of the jamming transition Song et al. (2008); Ramaswamy et al. (2023).
@@ Methods.
Shear protocol. Particles are randomly distributed at an initial volume fraction $\varphi_{0}=0.02$. The system is then rapidly quench compressed to the target $\varphi$. The CAQS is performed under the Lees-Edwards boundary conditions Lees and Edwards (1972), with a fixed $\varphi$. The strain $\gamma$ is varied stepwise, rather than continuously, between 0 and $\gamma_{\rm max}$. We use a strain step $\delta\gamma=0.1$ to generate the phase diagram in Fig. 1a, and $\delta\gamma=0.01$ for other results. At each step, particle positions are shifted according to $x_{i}\rightarrow x_{i}+\delta\gamma y_{i}$, and then the system’s energy is minimized using the FIRE algorithm Bitzek et al. (2006). During a cycle, the strain is varied as $\gamma=0\rightarrow\gamma_{\rm max}\rightarrow 0$. We set $\gamma_{\rm max}=0.7$ in the main text and $\gamma_{\rm max}=1$ in the SI for the 2D model, and $\gamma_{\rm max}=0.5$ for the 3D model (SI). The number of cycles is represented by $t$ with unit oscillation period ($t=1$). In 2D, the maximum number of cycles is $t_{\rm max}=250$ for $0.839\leq\varphi\leq{0.849}$ and $t_{\rm max}=4000$ for $0.839{0.833}\leq\varphi<$ 0.839, while in 3D, $t_{\rm max}=144$ for $2000$, and $t_{\rm max}=50$ for $N=512$ and $1000$. We generate $N_{\rm s}=4000-12000$ independent samples at each $\varphi$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ configurations are collected to compute statistical quantities.
@@ Acknowledgements.
We warmly thank Hajime Yoshino, Stephen Teitel, Bulbul Chakraborty, Olivier Dauchot, Gilles Tarjus for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
Supplementary Information
Figure S1: Additional data on the distribution $p(Z)$ in two dimensions. (a) $p(Z)$ for a few different $f_{\rm th}$ ($\varphi=0.841$ and $N=1000$). (b) Full probability distribution $\tilde{p}(Z)$ at a few different $\varphi$ near $\varphi_{\rm J}^{\infty}=0.8432$ ($N=1000$ and $f_{\rm th}=10^{-14}$).
Figure S2: Additional data on the jamming fraction $F_{\rm J}$ in two dimensions. (a) $F_{\rm J}$ as a function of $\varphi$, for different strain steps $\delta\gamma$, with fixed $\gamma_{\rm max}=0.7$ and $N=512$. (b) $F_{\rm J}$ as a function of $\varphi$, for different maximum strains $\gamma_{\rm max}$, with fixed $\delta\gamma=10^{-2}$ and $N=512$. (c) Data collapse of $F_{\rm J}(\varphi,N)$ as a function of rescaled variable $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, with $\varphi_{\rm J}^{\infty}=0.8432$, $\delta\gamma=10^{-1}$ and $\gamma_{\rm max}=0.7$.
Figure S3: Distribution of the coordination number in two dimensions for $\gamma_{\rm max}=1$. (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8435$. (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents the fitting according to Eq. (4),with two fitting parameters $u=0.051$ and $\sigma_{\varphi}=0.043$.
Figure S4: Susceptibilities in two dimensions for $\gamma_{\rm max}=1$. (a) The disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$ are plotted as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid black lines in (a-c) are Eqs. (5), (6) and (7) respectively. The dashed red line in (c) is the constant term in Eq. (7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$. To draw these lines, we use $u=0.051$ and $\sigma_{\varphi}=0.043$ obtained from the fitting in Fig. S3b. Symbols in (a-c) have the same meanings as indicated in the legend of (a).
Figure S5: Distribution of the coordination number in three dimensions. (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.648$. (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm j}^{\infty})N^{1/2}$. The solid line represents the fitting according to Eq. (4), with two fitting parameters $u=0.077$ and $\sigma_{\varphi}=0.051$.
Figure S6: Susceptibilities in three dimensions. (a) The disconnected susceptibility $\chi_{\rm dis}$ rescaled by $N$, (b) the connected susceptibility $\chi_{\rm con}$ rescaled by $N^{1/2}$, and (c) the ratio between $\chi_{\rm dis}$ and $\chi_{\rm con}^{2}$ are plotted as functions of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid black lines in (a-c) are Eqs. (5), (6) and (7) respectively. The dashed red line in (c) is the constant term in Eq. (7), i.e., $\chi_{\rm dis}/\chi_{\rm con}^{2}=\pi\sigma_{\varphi}^{2}/2$. To draw these lines, we use $u=0.077$ and $\sigma_{\varphi}=0.051$ obtained from the fitting in Fig. S5b. Symbols in (a-c) have the same meanings as indicated in the legend of (a).